The length of a rectangle is 2 in. more than twice its width. If the perimeter of the rectangle is 28 in., find the width of the rectangle.
Answer by Cromlix(4381)

Width = x Length = (2 + 2x)
Perimeter = 2 * length + 2 + width
= 2(2 + 2x) + 2x
= 4 + 4x + 2x
= 4 + 6x
Perimeter = 28 ins.
4 + 6x = 28 ins.
6x = 28 - 4 ins.
6x = 24 ins
x = 4 ins
Therefore width = 4 ins.
length= 10ins